OpenStudy (anonymous):

Let f(x)=log4x and g(x)=4^x x= 2

OpenStudy (anonymous):

\[f(x)=\log_4(x), g(x)=4^x\] they are inverse functions what does \(x=2\) have to do with it?

OpenStudy (anonymous):

Because I'm evaluating some equations and I know how to do the equations i just can`t figure out the part in the question

OpenStudy (anonymous):

evaluate \[\log_4(2)\]?

OpenStudy (anonymous):

same as solving \[4^x=2\] and since \[\sqrt4=4^{\frac{1}{2}}=2\] you have \[\log_4(2)=\frac{1}{2}\]

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

yw